OpenStudy (anonymous):

The midpoint of line UV is (5,-11). The coordinates of one endpoint are U(3,5). Find the coordinates of endpoint V.

OpenStudy (campbell_st):

let the other end point be V(x, y) the midpoint formula is basically add the 2 values then divide the answer by 2 so you have the answer.... so for x: 5 = (3 + x)/2 solve for x for y: -11 = (5 + y)/2 solve for y the answers will be the ordered pair V

OpenStudy (anonymous):

Okay, still don't know how to do that

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well for each if you multiply both sides of the equation by 2 5*2 = (3 + x)/2*2 which means you get 10 = 3 + x solve for y same process for y

OpenStudy (campbell_st):

oops should read solve for x... same process for finding y